40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1130     Entry time: Tue Mar 26 22:42:23 2013
Author: Evan 
Type: DailyProgress 
Category: optic 
Subject: Mode matching extravaganza 

Tara and I repositioned the QWP and PBS immediately preceding the periscope so that we could move the 64.4-mm ROC lens closer to the cavity. For space reasons, this lens is now forked directly to the table rather than mounted on a translation stage. I tried for a while to adjust this and the 38.6-mm ROC lens to improve the mode matching, but I can't seem to do much better than 80% visibility. We may have to adjust the 103-mm ROC lens directly after the PMC in order to go further.

In better news, we were able to couple some power into the fiber that runs into the ATF. The beam is picked off with the PBS immediately following the EAOM and then sent through two mode-matching lenses and a HWP before hitting the fiber. We're sending 10 mW in and currently getting 0.85 mW out. More work is needed to get the polarization correct and to improve the coupling efficiency. This setup will probably have to be redone at some point, since the current pickoff beam is downstream of the cavity EOM and therefore has sidebands on it. Also, we will have to redo the coupling if we touch the 103-mm ROC lens to improve the cavity mode matching.

ELOG V3.1.3-